9/11 Pager Messages
		
			2001-09-11
			06:37:41
			Arch
			[0509613]
			D
			ALPHA
			 (11)Initial X4236083 05:13 09/11/2001. PROB: Funlove Virus at KCMART IMP: Affecting 30 workstation and Possibly Peoplesoft STATUS: Bridge is in progress with Sitman(Brad) 800-733-4822*3. NXTUPD: 07:30 9/11/2001...JK/CR/emt
		 
		
			2001-09-11
			08:45:59
			Arch
			[0509613]
			D
			ALPHA
			 (12)Update X4236083. PROB: Funlove Virus at KCMART IMP: Affecting 33 workstations. STATUS: Desktop technicians are on site and addressing individual workstations. No servers infected at this time. Peoplesoft has been checked and cleared. Bridge
		 
		
			2001-09-11
			08:46:01
			Arch
			[0509613]
			D
			ALPHA
			 (13)is in progress with Sitman(Anita) 800-733-4822*3. NXTUPD: 09:37 9/11/2001...Chris/EM
		 
		
			2001-09-11
			09:07:55
			Arch
			[0509613]
			D
			ALPHA
			 (14)Initial 4236478 08:04 09/11/2001. PROB: Zephyr Hills unable to access CCPS.  System timing out. IMP: Customer Reps are unable to process
		 
		
			2001-09-11
			09:07:57
			Arch
			[0509613]
			D
			ALPHA
			  credit checks. STATUS: Conference call initiated to address issue. Babette/SitMan  800-733-4822*3. NXTU
		 
		
			2001-09-11
			09:08:23
			Arch
			[0509613]
			D
			ALPHA
			 (15)PD: 10:06 9/11/2001...Dave/BD/EM
		 
		
			2001-09-11
			10:07:23
			Arch
			[0509613]
			D
			ALPHA
			 (16)Restoral 4236478. PROB: Zephyr Hills unable to access CCPS.  System timing out. IMP: Customer Reps are unable to process credit checks. STATUS: Instance on the server was bounced to resolve issue. Verified resto
		 
		
			2001-09-11
			10:07:25
			Arch
			[0509613]
			D
			ALPHA
			 red 08:37 CDT Babette/SitMan
		 
		
			2001-09-11
			10:07:55
			Arch
			[0509613]
			D
			ALPHA
			 (17)800-733-4822*3....Dave/BD/EM
		 
		
			2001-09-11
			11:00:45
			Arch
			[0509613]
			D
			ALPHA
			 (18)ROC Event Initial SEV 2 PROBLEM: Due to overwhelming traffic to Operator Service.   JVL/OPS has been busied out as well as 1 span of every ITTO TG in PHX and LEN. IMPACT: Actions have been performed
		 
		
			2001-09-11
			11:00:53
			Arch
			[0509613]
			D
			ALPHA
			 (19)in response to excessive traffic to operator services.  Call completion time has been delayed.   STATUS: We are controlling traffic into the operator services platform. ETR: UNKNOWN TICKET: None /ESOC
		 
		
			2001-09-11
			11:47:39
			Arch
			[0509613]
			D
			ALPHA
			 (20)ROC Event Update SEV 2 PROBLEM:
		 
		
			2001-09-11
			11:47:41
			Arch
			[0509613]
			D
			ALPHA
			  Due to overwhelming traffic to the Operator Services platform.  ESOCC has busied out all ISDN traffic to JVL and 1span of every ITTO TG in LEN and PHX. IMPACT: Call
		 
		
			2001-09-11
			11:47:43
			Arch
			[0509613]
			D
			ALPHA
			 (21)completion time has been delayed. STATUS: Traffic is being routed to PHX and LEN OPS. ETR: unknown TICKET: UNKNOWN /ESOCC
		 
		
			2001-09-11
			11:57:49
			Arch
			[0509613]
			D
			ALPHA
			 (22)Restoral X4236083. PROB: Funlove Virus at KCMART IMP: Affecting 33 workstati
		 
		
			2001-09-11
			11:57:51
			Arch
			[0509613]
			D
			ALPHA
			 ons. STATUS: Restoral 10:15cdt. Only 5 machines were determined to be infected. Machines have been cleaned....Brandi/EM
		 
		
			2001-09-11
			12:10:37
			Arch
			[0509613]
			D
			ALPHA
			 (23)FYI...........MD Relay has declared a state of emergency.  The center will close at 1130cst and will reopen tomorrow 9/12 at 0530cst.
		 
		
			2001-09-11
			12:31:58
			Arch
			[0509613]
			D
			ALPHA
			 (24)ROC Event Update SEV 2 PROBLEM: Due to overwhelming traffic to the Operator Services platform.  ESOCC has busied out all ISDN traffic to JVL and 1sp
		 
		
			2001-09-11
			12:32:00
			Arch
			[0509613]
			D
			ALPHA
			 an of every ITTO TG in LEN and PHX. IMPACT: Call
		 
		
			2001-09-11
			12:32:02
			Arch
			[0509613]
			D
			ALPHA
			 (25)completion time has been delayed. STATUS: Traffic is being routed to PHX and LEN OPS. ETR: unknown TICKET: UNKNOWN /ESOCC
		 
		
			2001-09-11
			13:20:39
			Arch
			[0509613]
			D
			ALPHA
			 (26)Update NONE. Chicago Receivables office has been closed at 12:00 CDT.  Calls
		 
		
			2001-09-11
			13:20:41
			Arch
			[0509613]
			D
			ALPHA
			  have been rerouted to SAC2.  Chicago will reopen at 07:00 CDT on Wednesday 9/12/01
		 
		
			2001-09-11
			13:21:11
			Arch
			[0509613]
			D
			ALPHA
			 (27)Update NONE. Chicago Receivables office has been closed at 12:00 CDT.  Calls have been rerouted to SAC2.  Chicago will reopen at 07:00 CDT on Wednesday 9/12/01
		 
		
			2001-09-11
			13:36:24
			Arch
			[0509613]
			D
			ALPHA
			 (28)Initial X4229941 12:30 09/11/2001. PROB: WebOE - Application errors in Short and Long Forms. IMP: Customers are unable to order new service or perform maintenance on their accounts.
		 
		
			2001-09-11
			13:36:26
			Arch
			[0509613]
			D
			ALPHA
			  STATUS: Bridge initiated with Sitman/Babette 800-733-4822*
		 
		
			2001-09-11
			13:36:59
			Arch
			[0509613]
			D
			ALPHA
			 (29)3. NXTUPD: 14:30 9/11/2001...Ryan/jb/em
		 
		
			2001-09-11
			13:40:42
			Arch
			[0509613]
			D
			ALPHA
			 (30)Update 4232857. PROB: IEX Time stamp issue. IEX f
		 
		
			2001-09-11
			13:40:44
			Arch
			[0509613]
			D
			ALPHA
			 eeding incorrect time data to Aspect switches. Causing incorrect time stamps. IMP: Totalview losing data. Clients are unable to pull reports for their customer reps. STATUS: Aspect has made t
		 
		
			2001-09-11
			13:41:22
			Arch
			[0509613]
			D
			ALPHA
			 (31)he changes on Rel. 6 switches and are beginning the
		 
		
			2001-09-11
			13:41:24
			Arch
			[0509613]
			D
			ALPHA
			 process of processing the reports.  NXTUPD: 15:00 9/11/2001...Brandi/EM
		 
		
			2001-09-11
			14:29:22
			Arch
			[0509613]
			D
			ALPHA
			 (32)ROC Event Final SEV 2 PROBLEM: Due to overwhelming traffic to the Operator Services platform ESOCC busied ISDN to JVL, PHX and LEN.  All traffic has been restored to OPS at this time. SI: 0911ROC
		 
		
			2001-09-11
			14:29:23
			Arch
			[0509613]
			D
			ALPHA
			 (1)TICKET: UNKNOWN /ESOCC
		 
		
			2001-09-11
			15:38:19
			Arch
			[0509613]
			D
			ALPHA
			 (2)Re
		 
		
			2001-09-11
			15:38:21
			Arch
			[0509613]
			D
			ALPHA
			 storal X4229941. PROB: Web OE - App errors in Short and Long Forms. IMP: Customers cannot order new service or perform maintenance on accounts. STATUS: Corrupted data in Sales Person tables corrected. RESTORED at 14:20 CT... Scott/DAS
		 
		
			2001-09-11
			17:11:00
			Arch
			[0509613]
			D
			ALPHA
			 (3)Update 4232857. PROB: IEX Time stamp issue. IEX feeding incorrect time data to Aspect switches. Causing incorrect time stamps. IMP: Totalview losing data. Clients are unable to pull reports for the
		 
		
			2001-09-11
			17:11:02
			Arch
			[0509613]
			D
			ALPHA
			 ir customer reps. STATUS: Problem has been
		 
		
			2001-09-11
			17:11:30
			Arch
			[0509613]
			D
			ALPHA
			 (4)corrected.  Currently working on restoring lost data. NXTUPD: 18:19 9/11/2001...Peasley/em
		 
		
			2001-09-11
			17:15:15
			Arch
			[0509613]
			D
			ALPHA
			 (5)Traffic from Sac4 has been rerouted. Sac4 is closed for this evening and will resume operation tommorrow morning.
		 
		
			2001-09-11
			18:03:32
			Arch
			[0509613]
			D
			ALPHA
			 (6)Update 4232857. PROB: IEX Time stamp issue. IEX feeding incorrect time data to Aspect switches. Causing incorrect time stamps. IMP: Totalview losing data. Clients are unable to pull reports for their customer reps. STATUS: Prob
		 
		
			2001-09-11
			18:03:33
			Arch
			[0509613]
			D
			ALPHA
			 lem has been
		 
		
			2001-09-11
			18:04:02
			Arch
			[0509613]
			D
			ALPHA
			 (7)corrected.  All centers on the Aspect Rel.6 are now updating, and we are still waiting for fixes for the Rel.7s that IEX changed by mistake. NXTUPD: 19:00 9/11/2
		 
		
			2001-09-11
			18:04:04
			Arch
			[0509613]
			D
			ALPHA
			 001...Brandi/EM
		 
		
			2001-09-11
			19:49:40
			Arch
			[0509613]
			D
			ALPHA
			 (8)Restoral 4232857. PROB:
		 
		
			2001-09-11
			19:49:42
			Arch
			[0509613]
			D
			ALPHA
			  IEX Time stamp issue. IEX feeding incorrect time data to Aspect switches. Causing incorrect time stamps. IMP: Totalview losing data. Clients are unable to pull reports for their customer reps. STATUS: Problem has bee
		 
		
			2001-09-11
			19:50:09
			Arch
			[0509613]
			D
			ALPHA
			 (9)n corrected.  All centers on the Aspect Rel.6 and 7 switches have been processed resolving @ 18:37cdt....JoeEM